OpenStudy (anonymous):

7/6 divided by 9/3 times 27/4 divided by 3/2

OpenStudy (studygurl14):

When you divide by a fraction, multiply by its reciprocal. So you get: \(\Large \frac{7}{6}\times\frac{3}{9}\times\frac{4}{27}\times\frac{2}{3}\)